# Unveiling Entity-Level Unlearning for Large Language Models: A Comprehensive Analysis

### v3

@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Meta, 2024) pre-trained on extensive corpora have achieved significant success in various downstream tasks (Kamalloo et al., 2023; Seegmiller et al., 2024). However, training data often contains undesirable information, such as toxic texts (Lu et al., 2022), privacy content (Liu et al., 2024a) and copyrighted information (Karamolegkou et al., 2023). These issues raise security and legal concerns, hindering the practical application of LLMs (Yao et al., 2024; Das et al., 2024). To tackle this, Machine Unlearning (Zhang et al., 2023; Lu et al., 2024; Bhardwaj et al., 2024) has gradually been applied to LLMs due to its effectiveness and cost-efficiency. These refined techniques, now known as LLM Unlearning (Yao et al., 2023; Liu et al., 2024b, c), have become a mainstream approach for removing undesirable knowledge from the model by applying post-hoc modifications to target models.
@@
Towards this direction, the current unlearning paradigm involves applying an algorithm on a forget set which contains the undesirable knowledge (Zhang et al., 2024; Maini et al., 2024). However, most existing research focuses on Instance-level Unlearning tasks, which address isolated sensitive content (Li et al., 2024; Ji et al., 2024), while overlooking the deletion of entire entities, which is crucial in many real-world scenarios, such as removing ‘Harry Potter’ for copyright protection Eldan and Russinovich (2024). To address this gap, we formally define a novel task named Entity-level Unlearning. As illustrated in Figure 1, this task represents a significant divergence from instance-level unlearning due to the different nature of the information being targeted for removal. Specifically, instance-level unlearning focuses on removing specific facts, which are predefined and can be directly erased. In contrast, entity-level unlearning addresses the removal of an entire entity, requiring the deletion of all knowledge associated with the entity within the model, which is referred to as the target set. However, in real-world scenarios, the knowledge about an entity in the model is an abstract set that cannot be predefined. Consequently, entity-level unlearning requires an additional step to construct a specific forget set containing the entity-related knowledge for deletion.
@@ 3.1 Experimental Setup
  • Gradient Ascent (Grad. Ascent) (Yao et al., 2023), which is one of the most straightforward unlearning methods, reduces the likelihood of original answers on the forget set $S_{F}$.
@@
  • Negative Preference Optimization is an efficient unlearning method that requires only providing a negative response during preference optimization. We adopt the algorithm with the same restriction on the retain set $S_{R}$ as Gradient Difference (NPO-GD), which has been proven to outperform other variants (Zhang et al., 2024).
@@ 4.1 Effect of Forget Set
An intuitive approach to increase knowledge coverage is by increasing the size of the constructed forget set. Therefore, we further explore the performance of the algorithms on constructed forget sets across different sizes. Specifically, we expand the constructed forget set size for each entity and evaluate the performance of the unlearning algorithms across five different sizes. As illustrated in Figure 3, the forget quality of the five unlearning algorithms gradually improves as the size of the constructed forget set increases. However, except for the Pref. Opt. method, the model utility of the unlearned models produced by the other algorithms demonstrates a noticeable decline. This decline in model utility, when compared to the trends in Figure 2, is likely due to the larger number of unlearning steps required as the constructed forget set size grows. These findings suggest that, for most unlearning algorithms, expanding the constructed forget set size leads to a trade-off between forget quality and model utility. Furthermore, according to section 3.2, although the Pref. Opt. method maintains a high generation probability for the forget knowledge, it fails to achieve true deletion, as the forgotten information can still be easily retrieved through fill-in-the-blank questions (Wang et al., 2024).
@@ 5.1 Algorithms of LLM Unlearning
Current unlearning algorithms for LLMs (Cao and Yang, 2015; Yu et al., 2023; Jang et al., 2022) focus on minimizing the impact of undesirable training data while preserving the integrity of other model knowledge based on the Forget Set and Retain Set. Specifically, these methods (Jang et al., 2022; Chen and Yang, 2023) apply Gradient Ascent (GA) on the Forget Set to unlearn and add additional auxiliary loss on the Retain Set akin to gradient descent (Liu et al., 2022) and KL minimization (Maini et al., 2024) to mitigate undesired effects. Additionally, inspired by the alignment capabilities of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022), several researchers adopt preference optimization methods to refine model outputs, like Direct Policy Optimization (DPO) (Rafailov et al., 2024) and Distributional Dispreference Optimization (D2O) (Duan et al., 2024). Moreover, some researchers utilize model editing methods to remove sensitive information by adjusting knowledge-related parameters (Wu et al., 2023). However, it has been proven that the deleted content can be reverse-engineered from the edited model (Patil et al., 2023).
@@ 5.2 Evaluations of LLM Unlearning
Recent research has introduced several benchmarks and tasks for LLM unlearning from various aspects (Ji et al., 2024; Li et al., 2024; Maini et al., 2024; Lynch et al., 2024; Jang et al., 2022). Among these, the Weapons of Mass Destruction Proxy (WMDP) benchmark (Li et al., 2024) specifically targets dangerous knowledge in biosecurity, cybersecurity, and chemical security. While these task settings primarily address the forgetting of explicit instances, they pay less attention to entity-level unlearning, which involves completely forgetting an entity. Notably, Eldan and Russinovich (2024) explored a particular task to unlearn the entity "Harry Potter", but it is proven that the knowledge is not entirely erased from the unlearned model (Shi et al., 2023). Maini et al. (2024) presented TOFU consisting of 200 fictitious author profiles to assess unlearning methods from Model Utility and Forget Quality. However, this benchmark only focused on ideal scenarios with exact forget sets. Similarly, the RWKU benchmark (Jin et al., 2024) chooses 200 real-world famous people as entities for unlearning, with a more practical task setting. Nonetheless, it still focuses on specific pieces of knowledge related to target entities rather than the entire entity, making it closer to instance-level unlearning.
@@ B.3 Details for Baselines
We evaluate five common baselines for the entity-level unlearning task, following Maini et al. (2024); Zhang et al. (2024). The specific approaches are as follows:
  • Gradient Ascent (Yao et al., 2023), one of the most straightforward and basic unlearning algorithms, updates the target model, which is parametrized by $\theta_{t}$ by maximizing the cross-entropy loss $\ell(x,\theta_{t})$ over the forget set $S_{F}$:
@@
    Where $L_{\text{NPO}}(\cdot)$ represents the loss function of basic NPO (Zhang et al., 2024);

### v6

@@ 1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023a, b; Meta, 2024) pre-trained on extensive corpora have achieved significant success in various downstream tasks (Kamalloo et al., 2023; Seegmiller et al., 2024). However, training data often contains undesirable information, such as toxic texts (Lu et al., 2022), privacy content (Liu et al., 2024a) and copyrighted information (Karamolegkou et al., 2023). These issues raise security and legal concerns, hindering the practical application of LLMs (Yao et al., 2024; Das et al., 2024). To tackle this, Machine Unlearning (Zhang et al., 2023; Lu et al., 2024; Bhardwaj et al., 2024) has gradually been applied to LLMs due to its effectiveness and cost-efficiency. These refined techniques, now known as LLM Unlearning (Yao et al., 2023a; Liu et al., 2024b, c), have become a mainstream approach for removing undesirable knowledge from the model by applying post-hoc modifications to target models.
@@
Towards this direction, the current unlearning paradigm involves applying an algorithm on a forget set which contains the undesirable knowledge (Zhang et al., 2024b; Maini et al., 2024). However, most existing research focuses on Instance-level Unlearning tasks, which address isolated sensitive content (Li et al., 2024; Ji et al., 2024), while overlooking the deletion of entire entities, which is crucial in many real-world scenarios, such as removing ‘Harry Potter’ for copyright protection Eldan and Russinovich (2024). To address this gap, we formally define a novel task named Entity-level Unlearning. As illustrated in Figure 1, this task represents a significant divergence from instance-level unlearning due to the different nature of the information being targeted for removal. Specifically, instance-level unlearning focuses on removing specific facts, which are predefined and can be directly erased. In contrast, entity-level unlearning addresses the removal of an entire entity, requiring the deletion of all knowledge associated with the entity within the model, which is referred to as the target set. However, in real-world scenarios, the knowledge about an entity in the model is an abstract set that cannot be predefined. Consequently, entity-level unlearning requires an additional step to construct a specific forget set containing the entity-related knowledge for deletion.
@@ 3.1 Experimental Setup
  • Gradient Ascent (Grad. Ascent) (Yao et al., 2023a), which is one of the most straightforward unlearning methods, reduces the likelihood of original answers on the forget set $S_{F}$.
@@
  • Negative Preference Optimization is an efficient unlearning method that requires only providing a negative response during preference optimization. We adopt the algorithm with the same restriction on the retain set $S_{R}$ as Gradient Difference (NPO-GD), which has been proven to outperform other variants (Zhang et al., 2024b).
@@ 4.1 Effect of Forget Set
An intuitive approach to increase knowledge coverage is by increasing the size of the constructed forget set. Therefore, we further explore the performance of the algorithms on constructed forget sets across different sizes. Specifically, we expand the constructed forget set size for each entity and evaluate the performance of the unlearning algorithms across five different sizes. As illustrated in Figure 3, the forget quality of the five unlearning algorithms gradually improves as the size of the constructed forget set increases. However, except for the Pref. Opt. method, the model utility of the unlearned models produced by the other algorithms demonstrates a noticeable decline. This decline in model utility, when compared to the trends in Figure 2, is likely due to the larger number of unlearning steps required as the constructed forget set size grows. These findings suggest that, for most unlearning algorithms, expanding the constructed forget set size leads to a trade-off between forget quality and model utility. Furthermore, according to section 3.2, although the Pref. Opt. method maintains a high generation probability for the forget knowledge, it fails to achieve true deletion, as the forgotten information can still be easily retrieved through fill-in-the-blank questions (Wang et al., 2024a).
@@ 5.1 Algorithms of LLM Unlearning
LLM unlearning (Liu et al., 2024b) has attracted rising attention owing to its potential to address privacy (Jang et al., 2022) and security (Barrett et al., 2023) concerns. In addition, the LLM unlearning method can effectively mitigate the hallucination problem (Huang et al., 2023) in LLMs and analyze the influence of the training knowledge (Zhao et al., 2024). Current unlearning algorithms for LLMs (Cao and Yang, 2015; Yu et al., 2023; Jang et al., 2022) focus on minimizing the impact of undesirable training data while preserving the integrity of other model knowledge based on the Forget Set and Retain Set. Specifically, these methods (Jang et al., 2022; Chen and Yang, 2023) apply Gradient Ascent (GA) on the Forget Set to unlearn and add additional auxiliary loss on the Retain Set akin to gradient descent (Liu et al., 2022) and KL minimization (Maini et al., 2024) to mitigate undesired effects. Additionally, inspired by the alignment capabilities of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022), several researchers adopt preference optimization methods to refine model outputs, like Direct Policy Optimization (DPO) (Rafailov et al., 2024) and Distributional Dispreference Optimization (D2O) (Duan et al., 2024). Moreover, some researchers utilize model editing methods (Yao et al., 2023b; Feng et al., 2023; Zhang et al., 2024a) to remove sensitive information by adjusting knowledge-related parameters (Wu et al., 2023; Wang et al., 2024b). However, it has been proven that the deleted content can be reverse-engineered from the edited model (Patil et al., 2023).
@@ 5.2 Evaluations of LLM Unlearning
Recent research has introduced several benchmarks and tasks for LLM unlearning from various aspects (Ji et al., 2024; Li et al., 2024; Maini et al., 2024; Lynch et al., 2024; Jang et al., 2022; tian2024forget). Among these, the Weapons of Mass Destruction Proxy (WMDP) benchmark (Li et al., 2024) specifically targets dangerous knowledge in biosecurity, cybersecurity, and chemical security. While these task settings primarily address the forgetting of explicit instances, they pay less attention to entity-level unlearning, which involves completely forgetting an entity. Notably, Eldan and Russinovich (2024) explored a particular task to unlearn the entity "Harry Potter", but it is proven that the knowledge is not entirely erased from the unlearned model (Shi et al., 2023). Maini et al. (2024) presented TOFU consisting of 200 fictitious author profiles to assess unlearning methods from Model Utility and Forget Quality. However, this benchmark only focused on ideal scenarios with exact forget sets. Similarly, the RWKU benchmark (Jin et al., 2024) chooses 200 real-world famous people as entities for unlearning, with a more practical task setting. Nonetheless, it still focuses on specific pieces of knowledge related to target entities rather than the entire entity, making it closer to instance-level unlearning.
@@ 6 Conclusion
Acknowledgements
Xiaocheng Feng is the corresponding author of this work. We thank the anonymous reviewers for their insightful comments. This work was supported by the National Natural Science Foundation of China (NSFC) (grant 62276078, U22B2059), the Key R&D Program of Heilongjiang via grant 2022ZX01A32, the International Cooperation Project of PCL, PCL2022D01 and the Fundamental Research Funds for the Central Universities (Grant No.HIT.OCEF.2023018). We also thank Du Xiaoman (Beijing) Science Technology Co., Ltd for supporting part of the computing resources and funding.
@@ B.3 Details for Baselines
We evaluate five common baselines for the entity-level unlearning task, following Maini et al. (2024); Zhang et al. (2024b). The specific approaches are as follows:
  • Gradient Ascent (Yao et al., 2023a), one of the most straightforward and basic unlearning algorithms, updates the target model, which is parametrized by $\theta_{t}$ by maximizing the cross-entropy loss $\ell(x,\theta_{t})$ over the forget set $S_{F}$:
@@
    Where $L_{\text{NPO}}(\cdot)$ represents the loss function of basic NPO (Zhang et al., 2024b);